OpenStudy (anonymous):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book? Enter your answer in the box.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Loser66

OpenStudy (itsbribro):

$16.22 i believe

OpenStudy (anonymous):

ok :)

OpenStudy (michele_laino):

after discount, the new price is: \[\Large 18 - \left( {\frac{{15}}{{100}} \cdot 18} \right) = ...?\]

OpenStudy (itsbribro):

oops i messed up lol

OpenStudy (anonymous):

153/10

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

is it 15.3 @Michele_Laino

OpenStudy (michele_laino):

please wait, I check your result

OpenStudy (anonymous):

ok :)

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

ok :)

OpenStudy (michele_laino):

now, the requested price is given by the subsequent computation: \[\Large 15.3 + \left( {\frac{6}{{100}} \cdot 15.3} \right) = ...?\]

OpenStudy (anonymous):

15.6

OpenStudy (michele_laino):

I got a different result

OpenStudy (anonymous):

what did you get?

OpenStudy (michele_laino):

hint: \[\Large \frac{6}{{100}} \cdot 15.3 = ...?\]

OpenStudy (anonymous):

0.918 @Michele_Laino

OpenStudy (michele_laino):

and 15.3+0.918=...?

OpenStudy (anonymous):

16.128

OpenStudy (anonymous):

i mean 16.218

OpenStudy (michele_laino):

yes! correct! That result can be rounded off to 16.22

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (michele_laino):

the final answer is 16.22

OpenStudy (anonymous):

are you sure?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok :)

OpenStudy (michele_laino):

here the rate of the car is: space/time, \[\Large \frac{{1/6}}{{3/5}} = \frac{1}{6} \cdot \frac{5}{3} = ...?\]

OpenStudy (michele_laino):

please express your result as a fraction

OpenStudy (anonymous):

1/10

OpenStudy (michele_laino):

I'm sorry, it is wrong

OpenStudy (anonymous):

1/10 is wrong?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

5/18?

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

that rate is the space traveled in 1 hour

OpenStudy (michele_laino):

so the requested distance is: distance= speed*time \[\Large {\text{distance}} = \frac{5}{{18}} \cdot 1 = ...?\]

OpenStudy (anonymous):

sorry i meant 5/18

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

is that the final answer?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

okay :)

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

part A) If I call with p the price in September, then I can write this equation: \[\Large \begin{gathered} p + \left( {\frac{{20}}{{100}} \cdot p} \right) = 1.5 \hfill \\ \hfill \\ or: \\ \\ 1.2p = 1.5 \hfill \\ \end{gathered} \] please solve for p

OpenStudy (anonymous):

p=1.25

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

is that the final answer for part A?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok now part b?

OpenStudy (michele_laino):

here, we note that: in september the earned money is: \[\Large 900 \cdot \left( {1.25 \cdot \frac{{40}}{{100}}} \right) = ...?\] whereas in october, the earned money is: \[\Large 700 \cdot \left( {1.5 \cdot \frac{{40}}{{100}}} \right) = ...?\] please compute both amounts of money

OpenStudy (anonymous):

i got 450 for september and 420 for october

OpenStudy (michele_laino):

correct! So, what is your answer?

OpenStudy (anonymous):

um its says choose one... for the first one these are the multiple choices.. september or october and the other choose one are... $30, $75, or $190

OpenStudy (michele_laino):

yes! 450-420=...?

OpenStudy (anonymous):

30

OpenStudy (michele_laino):

so what are your options?

OpenStudy (anonymous):

for which one... there are two answers for part b

OpenStudy (michele_laino):

yes! for part B) there are 2 options

OpenStudy (anonymous):

right

OpenStudy (michele_laino):

first option: is september or october?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

please choose

OpenStudy (anonymous):

september

OpenStudy (michele_laino):

correct! second option: is 30, 75 or 190?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

$30?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! so september and 30

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

part A) hint: your line passes at point (0,0), namely the origin

OpenStudy (anonymous):

i got... a,c and e

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thanks so much! :)

OpenStudy (michele_laino):

:)